Samsung has a correlation with the S&P 500 of 0.8. Samsung's standard deviation is 0.25 and the S&P 500 has a standard deviation of 0.22. The risk-free rate is 0.01, and the expected return on the S&P 500 is 0.05.

What is Samsung's expected return?

If you think the expected return is 12%, what is Samsung's alpha?
Answer #1

Beta = Correlation* Standard Deviation of Samsung/ Standard Deviation of S&P = 0.8*0.25/0.22 = 0.90901

Expected Return = Risk Free rate + Beta*(Market Return - Risk Free rate) = 0.01 + 0.8*0.25/0.22 *( 5%-1%) = 4.64%

If Expected return is 12% then alpha = 12% - 4.64% = 7.36%

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well